# FARCLUSS: Fuzzy Adaptive Rebalancing and Contrastive Uncertainty Learning for Semi-Supervised Semantic Segmentation

### v1

@@ Abstract
Semi-supervised semantic segmentation (SSSS) faces persistent challenges in effectively leveraging unlabeled data, such as ineffective utilization of pseudo-labels, exacerbation of class imbalance biases, and neglect of prediction uncertainty. Current approaches often discard uncertain regions through strict thresholding favouring dominant classes. To address these limitations, we introduce a holistic framework that transforms uncertainty into a learning asset through four principal components: (1) fuzzy pseudo-labeling, which preserves soft class distributions from top-K predictions to enrich supervision; (2) uncertainty-aware dynamic weighting, that modulate pixel-wise contributions via entropy-based reliability scores; (3) adaptive class rebalancing, which dynamically adjust losses to counteract long-tailed class distributions; and (4) lightweight contrastive regularization, that encourage compact and discriminative feature embeddings. Extensive experiments on benchmarks demonstrate that our method outperforms current state-of-the-art approaches, achieving significant improvements in the segmentation of under-represented classes and ambiguous regions.
@@
journal: Pattern Recognition
@@ 1 Introduction
Semantic segmentation, defined as the process of assigning class labels to every pixel in an image, supports a range of critical applications, including autonomous driving and medical diagnostics. By facilitating pixel-level scene analysis, it supports systems that demand fine-grained environmental understanding [5]. Nonetheless, fully supervised segmentation techniques usually depend on extensive datasets with detailed annotations; this makes the method costly in fields like urban scene analysis or healthcare, where manual labelling is both expensive and time-consuming [1]. To mitigate this dependency on labeled data, semi-supervised semantic segmentation (SSSS) has emerged as a promising alternative. By leveraging a modest set of labeled images alongside abundant unlabeled data, SSSS frameworks strive to achieve performance comparable to fully supervised models while substantially reducing annotation requirements. Early work in this area centered on pseudo-labeling [12], wherein a teacher model generates approximate labels for unlabeled data that a student model subsequently learns from. Later research introduced consistency regularization [13], enforcing prediction invariance under various geometric or photometric transformations. Although these techniques have shown improvements, they remain constrained by three principal issues: suboptimal use of pseudo-labels, class imbalance in long-tailed datasets, and inadequate handling of prediction uncertainty.
Figure 1: Overview of consistency-regularization-based semi-supervised semantic segmentation (SSSS) frameworks during training with unlabeled data. ps, pf, and psf are predictions of strong input perturbation stream, feature perturbation stream, and strong input + feature perturbation stream. $f(\theta)$ means exponential moving average (EMA) updated model.
@@
A closer examination reveals three principal shortcomings in current SSSS methods: ineffective utilization of pseudo-labels, class imbalance biases, and excessive computational cost. Although recent approaches attempt to address these issues, they often lack comprehensive solutions. Firstly, most methods treat pseudo-label uncertainty as a liability. They discard low-confidence predictions via strict thresholding, overlooking subtle probabilistic relationships among plausible classes. For instance, a pixel at the boundary of ’sidewalk’ and ’road’ may carry near-equal probabilities for both classes, yet is often removed, forfeiting valuable supervisory signals. UniMatch [3] combines feature- and output-level consistency but relies on fixed confidence thresholds to generate hard pseudo-labels. As a result, ambiguous predictions are eliminated, especially near object boundaries or occluded regions,leading to loss of information.
@@
Secondly, class imbalance remains inadequately addressed. Although some methods adjust global thresholds, they fail to rebalance losses based on per-batch class distributions, perpetuating bias toward majority classes. PS-MT [2] employs progressive self-training with momentum teachers and adaptive thresholds but lacks explicit mechanisms for class rebalancing. Consequently, dominant categories (e.g., "road" or "sky" in urban scenes) are over-represented in pseudo-labels.
@@
Thirdly, contrastive learning frameworks often incur high computational cost due to extensive feature alignments, limiting scalability on large datasets. ReCo [4] improves feature discriminability using region-level contrastive learning, but its pairwise embedding comparisons introduce considerable overhead. Cross-pseudo supervision (CPS) [6] trains dual networks that supervise each other, yet applies uniform weighting to pseudo-labels, neglecting the need to treat uncertain regions differently.
@@
In response, this work proposes a unified framework that systematically addresses common challenges in semi-supervised segmentation, including pseudo-label noise, class imbalance, and computational overhead (Fig. 1). First, we introduce fuzzy pseudo-labeling, which derives soft label distributions from the top-K class probabilities at each pixel. By preserving uncertainty (unlike previous methods’ strict thresholding), our method treats ambiguity as a constructive supervisory cue. This enables the model to learn from regions like unclear boundaries (e.g., pavement edges between "car" and "road") instead of discarding them.
@@
Third, we propose adaptive class rebalancing to mitigate class imbalance by dynamically adjusting loss weights according to batch-level pseudo-label frequencies. This ensures under-represented categories like "traffic sign" or "tumor" receive adequate optimization focus without manual tuning.
@@
While existing methods have tackled individual challenges such as ineffective pseudo-label utilization, class imbalance, and high computational cost, they often fall short of addressing these issues simultaneously within a unified and performant framework. The main contributions of this work are summarized as follows:
@@ 2.1 Teacher-Student Frameworks and Pseudo-Label Refinement
Semi-supervised semantic segmentation (SSSS) has progressed remarkably through teacher-student paradigms, wherein models generate pseudo-labels from unlabeled data to guide training. Early work such as FixMatch [7] employed confidence thresholding to filter pseudo-labels, leading to high-confidence yet inaccurate predictions reinforced model errors (confirmation bias). Cross Pseudo Supervision (CPS) [6] mitigated this issue by training two networks jointly in a supervised manner using each other’s pseudo-labels, thereby diversifying the learning signal. PseudoSeg [8] enhanced reliability through augmented prediction fusion, while Directional Context-Aware Consistency (DC-Loss) [9] encouraged alignment of class-specific features in differing contexts. More recent approaches, including the Error Localization Network (ELN) [10] and U²PL [11], introduced modules to detect, filter, or reinterpret uncertain predictions as negative supervision. However, most of these methods still rely on strict thresholding, discarding ambiguous pixels and thereby limiting data usage. By contrast, our framework employs fuzzy pseudo-labeling to retain the top-K class probabilities for each pixel, preserving nuanced inter-class relationships and providing richer supervision even in low-confidence regions.
@@ 2.3 Class Imbalance and Rare Class Enhancement
Class imbalance continues to pose significant challenges for SSSS, as pseudo-labels often exhibit a bias toward frequently occurring classes. DARS [18] aligned pseudo-label distributions with those from labeled data, and Adaptive Equalization Learning (AEL) [19] amplified underrepresented classes through adaptive loss weighting. Augmentation-based approaches, including Dynamic Cross-Set Copy-Paste (DCSCP) [20] and iMAS [21], artificially increased rare-class samples or dynamically adjusted supervision based on confidence. While these methods successfully address class imbalance, they typically rely on global class distributions, overlooking variations at the batch level. In contrast, our adaptive class rebalancing technique leverages per-batch pseudo-label frequencies to scale losses, ensuring that classes like traffic sign in Cityscapes receive increased focus when they appear infrequently. Moreover, we employ a prototype-based contrastive regularization strategy, wherein class centroids derived from high-confidence pseudo-labels enhance intra-class compactness. Unlike ReCo [4], which entails computationally expensive pairwise comparisons, this prototype-based approach offers improved feature separability for minority classes with minimal overhead.
@@ 2.4 Computational Efficiency and Lightweight Learning
Contemporary SSSS approaches frequently carry a high computational burden due to multi-model architectures or complex regularization. Cross-Consistency Training (CCT) [22] reduced this burden by training multiple decoder heads with a shared encoder, while UniMatch [3] illustrated that a single-network FixMatch [7] can achieve performance comparable to multi-model solutions. Multi-Granularity Distillation (MGD) [23] further improved efficiency via hierarchical knowledge distillation. Building on these principles, our framework avoids dual-network configurations like CPS [6] and auxiliary modules such as ELN [10]. Through single-pass fuzzy pseudo-labeling, entropy-based weighting, and prototype-based contrastive learning, we preserve scalability on large datasets without sacrificing segmentation accuracy.
@@ 3 Method
Our method integrates fuzzy pseudo-labeling, uncertainty-based dynamic weighting, adaptive class rebalancing, and efficient contrastive regularization to address inefficient pseudo-label utilization, class imbalance biases, and underutilization of uncertain predictions. An overview of our proposed method is shown in Fig. 3.
@@ 3.7.2 Practical Considerations
Algorithm 1 FARCLUSS pseudocode in PyTorch style
# f: encoder g + decoder h; f_ema: EMA teacher; aug_w/aug_s: perturbations
criterion_ce = nn.KLDivLoss(log_target=False, reduction="none")
criterion_con = nn.CosineSimilarity(dim=1)
m = 0.99 # EMA momentum # eq. 1
for x in loader_u:
# generate weak/strong views
xw = aug_w(x)
xs1, xs2 = aug_s(xw), aug_s(xw)
# teacher inference & fuzzy labels
with torch.no_grad():
pw = f_ema(xw)
tk = pw.topk(K, 1).indices
fuzzy = torch.zeros_like(pw).scatter_(1, tk, pw.gather(1, tk))
fuzzy = fuzzy / fuzzy.sum(dim=1, keepdim=True) #eq. 2
H = - (pw * pw.log()).sum(dim=1) / math.log(pw.shape[1])
w_px = 1 - H # pixel/instance weight # eq. 6
# student features & mask
feat = g(torch.cat([xs1, xs2])) # B*2, C, H, W
bs2, C, H, W = feat.shape
m0 = torch.bernoulli(0.5 * torch.ones(bs2//2, C, device=feat.device))
mask = torch.cat([m0, 1 - m0], 0).view(bs2, C, 1, 1)
pred = h(feat * mask)
# supervised loss on labeled batch
x_l, y_l = next(iter(loader_s))
loss_sup = nn.CrossEntropyLoss()(f(x_l).log_softmax(1), y_l)
# adaptive class rebalancing loss
lbl = fuzzy.argmax(1)
cnt = lbl.bincount(minlength=pw.size(1)).float()
w_cls = cnt.median() / (cnt + 1e-6) # eq. 7
loss_u_map = criterion_ce(pred.log_softmax(1), fuzzy) # B*2 × C × H × W
loss_u = (w_px.view(-1,1,1,1) * w_cls[lbl].view(-1,1,1,1) * loss_u_map).mean()
# prototype contrastive loss
sel = w_px > 0.5
fs = feat[sel] # N_sel, C, H, W
lb = lbl[sel]
fs_vec = fs.mean((2,3)) # N_sel × C
cnt_sel = lb.bincount(minlength=pw.size(1)).float()
proto = torch.zeros(pw.size(1), C, device=feat.device)
proto.index_add_(0, lb, fs_vec)
proto = proto / (cnt_sel.unsqueeze(1) + 1e-6) # eq. 9
loss_c = (1 - criterion_con(fs_vec, proto[lb])).mean() # eq. 10
# update student & EMA
loss = loss_sup + λ_u * loss_u + λ_c * loss_c # eq. 11
optimizer.zero_grad(); loss.backward(); optimizer.step()
for pt, ps in zip(f_ema.parameters(), f.parameters()):
pt.data.mul_(m).add_(ps.data, alpha=1-m)
@@ 4.1 Implementation Details
Our model is implemented using PyTorch and trained on Pascal VOC and Cityscapes datasets following common semi-supervised segmentation setups [3]. We build upon a DeepLabV3+ baseline architecture with ResNet-50 and ResNet-101 backbones [14, 15] with dilation rates $\{6,12,18\}$, and adopt standard mean-teacher EMA updates [7]. Training is performed for 80 epochs with a batch size of 8 for Pascal and 240 epochs for Cityscapes. The initial learning rate is set to $0.001$, following common practice.
@@ 4.2 Datasets and Semi-Supervised Setups
4.2.1 Pascal VOC (Classic & Blended
@@ 4.2.2 Cityscapes
The Cityscapes dataset [41] consists of 2,975 training images with fine-grained annotations and 500 validation images, featuring urban scenes labeled across 19 semantic classes. Despite its relatively smaller class space, Cityscapes offers high-resolution annotations (originally at 1024x2048 pixels), which include detailed labels for small and thin objects such as traffic lights. Model performance is evaluated by calculating the mIoU metric on the provided 500-image validation set.
@@ 4.2.3 Data Augmentation
Following common practice in semi-supervised segmentation [22, 19], we apply weak augmentation (e.g., random scaling, random horizontal flipping) for teacher inference and strong augmentation (e.g., random crops, color jitter, stronger geometric transformations) for student training. Weakly-augmented images yield more stable teacher predictions, while strongly-augmented images regularize the student to learn robust features under varied conditions.
@@ 4.3.3 Quantitative Analysis on Cityscapes
Figure 5: Qualitative results of our method, FARCLUSS with other state-of-the-art methods on the Cityscapes dataset. All methods are trained under the 1/8 (12%) setting for validation. Red rectangles highlight regions where segmentation performance is improved.

### v2

@@ Abstract
Semi-supervised semantic segmentation (SSSS) faces persistent challenges in effectively leveraging unlabeled data, such as ineffective utilization of pseudo-labels, exacerbation of class imbalance biases, and neglect of prediction uncertainty. Moreover, current approaches often discard uncertain regions through strict thresholding favouring dominant classes. To address these limitations, we introduce a holistic framework that transforms uncertainty into a learning asset through four principal components: (1) fuzzy pseudo-labeling, which preserves soft class distributions from top-K predictions to enrich supervision; (2) uncertainty-aware dynamic weighting, that modulate pixel-wise contributions via entropy-based reliability scores; (3) adaptive class rebalancing, which dynamically adjust losses to counteract long-tailed class distributions; and (4) lightweight contrastive regularization, that encourage compact and discriminative feature embeddings. Extensive experiments on benchmarks demonstrate that our method outperforms current state-of-the-art approaches, achieving significant improvements in the segmentation of under-represented classes and ambiguous regions.
@@
journal: Neural Networks
@@ 1 Introduction
Semantic segmentation, defined as the process of assigning class labels to every pixel in an image, supports a range of critical applications, including autonomous driving and medical diagnostics. By facilitating pixel-level scene analysis, it supports systems that demand fine-grained environmental understanding [5]. Nonetheless, fully supervised segmentation techniques usually depend on extensive datasets with detailed annotations; making it costly in fields like urban scene analysis or healthcare, where manual labelling is both expensive and time-consuming [1]. To mitigate this dependency on labeled data, semi-supervised semantic segmentation (SSSS) has emerged as a promising alternative. By leveraging a modest set of labeled images alongside abundant unlabeled data, SSSS frameworks strive to achieve performance comparable to fully supervised models while substantially reducing annotation requirements. Early work in this area centered on pseudo-labeling [12], wherein a teacher model generates approximate labels for unlabeled data that a student model subsequently learns from. Later research introduced consistency regularization [13], enforcing prediction invariance under various geometric or photometric transformations. Although these techniques have shown improvements, they remain constrained by three principal issues: suboptimal use of pseudo-labels, class imbalance in long-tailed datasets, and inadequate handling of prediction uncertainty.
@@
Although recent approaches attempt to address these issues, they often lack comprehensive solutions. Most methods treat pseudo-label uncertainty as a liability. They discard low-confidence predictions via strict thresholding, overlooking subtle probabilistic relationships among plausible classes. For instance, a pixel at the boundary of ’sidewalk’ and ’road’ may carry near-equal probabilities for both classes, yet is often removed, forfeiting valuable supervisory signals. UniMatch [3] combines feature- and output-level consistency but relies on fixed confidence thresholds to generate hard pseudo-labels. As a result, ambiguous predictions are eliminated, especially near object boundaries or occluded regions,leading to loss of information.
@@
Class imbalance remains inadequately addressed. Although some methods adjust global thresholds, they fail to rebalance losses based on per-batch class distributions, perpetuating bias toward majority classes. PS-MT [2] employs progressive self-training with momentum teachers and adaptive thresholds but lacks explicit mechanisms for class rebalancing. Consequently, dominant categories (e.g., "road" or "sky" in urban scenes) are over-represented in pseudo-labels.
@@
Methods that use contrastive learning to address previous mentioned limitations often incur high computational cost due to extensive feature alignments, limiting scalability on large datasets. ReCo [4] improves feature discriminability using region-level contrastive learning, but its pairwise embedding comparisons introduce considerable overhead. Cross-pseudo supervision (CPS) [6] trains dual networks that supervise each other, yet applies uniform weighting to pseudo-labels, neglecting the need to treat uncertain regions differently.
@@
This work proposes a unified framework that systematically addresses common challenges in semi-supervised segmentation, including pseudo-label noise, class imbalance, and computational overhead. First, we introduce fuzzy pseudo-labeling, which derives soft label distributions from the top-K class probabilities at each pixel. By preserving uncertainty (unlike previous methods’ strict thresholding), our method treats ambiguity as a constructive supervisory cue. This enables the model to learn from regions like unclear boundaries (e.g., pavement edges between "car" and "road") instead of discarding them.
@@
Third, we propose adaptive class rebalancing to mitigate class imbalance by dynamically adjusting loss weights according to batch-level pseudo-label frequencies. This ensures under-represented categories like "traffic sign" receive adequate optimization focus without manual tuning.
@@
While existing methods have tackled individual challenges such as ineffective pseudo-label utilization, class imbalance, and high computational cost, they often fall short of addressing these issues simultaneously within a unified framework. The main contributions of this work are summarized as follows:
@@ 2.1 Teacher-Student Frameworks and Pseudo-Label Refinement
Semi-supervised semantic segmentation (SSSS) has progressed remarkably through teacher-student paradigms, wherein models generate pseudo-labels from unlabeled data to guide training. Early work such as FixMatch [7] employed confidence thresholding to filter pseudo-labels, leading to high-confidence yet inaccurate predictions reinforced model errors (confirmation bias). Cross Pseudo Supervision (CPS) [6] mitigated this issue by training two networks jointly in a supervised manner using each other’s pseudo-labels, thereby diversifying the learning signal. PseudoSeg [8] enhanced reliability through augmented prediction fusion, while Directional Context-Aware Consistency (DC-Loss) [9] encouraged alignment of class-specific features in differing contexts. More recent approaches, including the Error Localization Network (ELN) [10] and $U^{2}\mathrm{PL}$ [11], introduced modules to detect, filter, or reinterpret uncertain predictions as negative supervision. However, most of these methods still rely on strict thresholding, discarding ambiguous pixels and thereby limiting data usage. By contrast, our framework employs fuzzy pseudo-labeling to retain the top-K class probabilities for each pixel, preserving nuanced inter-class relationships and providing richer supervision even in low-confidence regions.
@@ 2.3 Class Imbalance and Rare Class Enhancement
Class imbalance continues to pose significant challenges for SSSS, as pseudo-labels often exhibit a bias toward frequently occurring classes. DARS [18] aligned pseudo-label distributions with those from labeled data, and Adaptive Equalization Learning (AEL) [19] amplified underrepresented classes through adaptive loss weighting. Augmentation-based approaches, including Dynamic Cross-Set Copy-Paste (DCSCP) [20] and iMAS [21], artificially increased rare-class samples or dynamically adjusted supervision based on confidence. While these methods successfully address class imbalance, they typically rely on global class distributions, overlooking variations at the batch level. In contrast, our adaptive class rebalancing technique leverages per-batch pseudo-label frequencies to scale losses, ensuring that classes like traffic sign in Cityscapes receive increased focus when they appear infrequently.
@@ 2.4 Computational Efficiency and Lightweight Learning
Contemporary SSSS approaches frequently carry a high computational burden due to multi-model architectures or complex regularization. Cross-Consistency Training (CCT) [22] reduced this burden by training multiple decoder heads with a shared encoder, while UniMatch [3] illustrated that a single-network FixMatch [7] can achieve performance comparable to multi-model solutions. Multi-Granularity Distillation (MGD) [23] further improved efficiency via hierarchical knowledge distillation. Building on these principles, our framework avoids dual-network configurations like CPS [6] and auxiliary modules such as ELN [10]. Through single-pass fuzzy pseudo-labeling, entropy-based weighting, and prototype-based contrastive learning, wherein class centroids derived from high-confidence pseudo-labels enhance intra-class compactness, we preserve scalability on large datasets without sacrificing segmentation accuracy. Unlike ReCo [4], which entails computationally expensive pairwise comparisons, this prototype-based approach offers improved feature separability for minority classes with minimal overhead.
@@ 3 Method
Our method integrates fuzzy pseudo-labeling, uncertainty-based dynamic weighting, adaptive class rebalancing, and efficient contrastive regularization to address inefficient pseudo-label utilization, class imbalance biases, and underutilization of uncertain predictions. An overview of our proposed method is shown in Fig. 2 and the corresponding pseudocode is provided in Algorithm 1.
@@ 3.7.1 Combined Loss
Algorithm 1 FARCLUSS pseudocode in PyTorch style
# f: encoder g + decoder h; f_ema: EMA teacher; aug_w/aug_s: perturbations criterion_ce = nn.KLDivLoss(log_target=False, reduction="none") criterion_con = nn.CosineSimilarity(dim=1) m = 0.99 # EMA momentum # eq. 1 for x in loader_u: # generate weak/strong views xw = aug_w(x) xs1, xs2 = aug_s(xw), aug_s(xw) # teacher inference & fuzzy labels with torch.no_grad(): pw = f_ema(xw) tk = pw.topk(K, 1).indices fuzzy = torch.zeros_like(pw).scatter_(1, tk, pw.gather(1, tk)) fuzzy = fuzzy / fuzzy.sum(dim=1, keepdim=True) #eq. 2 H = - (pw * pw.log()).sum(dim=1) / math.log(pw.shape[1]) w_px = 1 - H # pixel/instance weight # eq. 6 # student features & mask feat = g(torch.cat([xs1, xs2])) # B*2, C, H, W bs2, C, H, W = feat.shape m0 = torch.bernoulli(0.5 * torch.ones(bs2//2, C, device=feat.device)) mask = torch.cat([m0, 1 - m0], 0).view(bs2, C, 1, 1) pred = h(feat * mask) # supervised loss on labeled batch x_l, y_l = next(iter(loader_s)) loss_sup = nn.CrossEntropyLoss()(f(x_l).log_softmax(1), y_l) # adaptive class rebalancing loss lbl = fuzzy.argmax(1) cnt = lbl.bincount(minlength=pw.size(1)).float() w_cls = cnt.median() / (cnt + 1e-6) # eq. 7 loss_u_map = criterion_ce(pred.log_softmax(1), fuzzy) # B*2 × C × H × W loss_u = (w_px.view(-1,1,1,1) * w_cls[lbl].view(-1,1,1,1) * loss_u_map).mean() # prototype contrastive loss sel = w_px > 0.5 fs = feat[sel] # N_sel, C, H, W lb = lbl[sel] fs_vec = fs.mean((2,3)) # N_sel × C cnt_sel = lb.bincount(minlength=pw.size(1)).float() proto = torch.zeros(pw.size(1), C, device=feat.device) proto.index_add_(0, lb, fs_vec) proto = proto / (cnt_sel.unsqueeze(1) + 1e-6) # eq. 9 loss_c = (1 - criterion_con(fs_vec, proto[lb])).mean() # eq. 10 # update student & EMA loss = loss_sup + λ_u * loss_u + λ_c * loss_c # eq. 11 optimizer.zero_grad(); loss.backward(); optimizer.step() for pt, ps in zip(f_ema.parameters(), f.parameters()): pt.data.mul_(m).add_(ps.data, alpha=1-m)
@@ 4.1 Implementation Details
Our model is implemented using PyTorch and trained on Pascal VOC and Cityscapes datasets following common semi-supervised segmentation setups [3]. We build upon a DeepLabV3+ baseline architecture with ResNet-50 and ResNet-101 backbones [14, 15] with dilation rates $\{6,12,18\}$, and adopt standard mean-teacher EMA updates [7]. Training is performed for 80 epochs for Pascal and 240 epochs for Cityscapes. The initial learning rate is set to $0.001$, following common practice.
@@ 4.2 Datasets and Semi-Supervised Setups
4.2.1 Pascal VOC (Classic & Blended)
@@ 4.2.2 Cityscapes
Cityscapes dataset [41] consists of 2,975 training images with fine-grained annotations and 500 validation images, featuring urban scenes labeled across 19 semantic classes. Despite its relatively smaller class space, Cityscapes offers high-resolution annotations (originally at 1024x2048 pixels), which include detailed labels for small and thin objects such as traffic lights. Model performance is evaluated by calculating the mIoU metric on the provided 500-image validation set.
@@ 4.2.3 Data Augmentation
Following common practice in semi-supervised segmentation [22, 19], we apply weak augmentation (e.g., random scaling, random horizontal flipping) for teacher inference and strong augmentation (e.g., random crops, color jitter, stronger geometric transformations) for student training. Weakly-augmented images yield more stable teacher predictions, while strongly-augmented images regularize the student to learn robust features under varied conditions [22].
@@ 4.3.4 Quantitative Analysis on Computational Cost
Figure 4: Qualitative results of our method, FARCLUSS with other state-of-the-art methods on the Cityscapes dataset. All methods are trained under the 1/8 (12.5%) setting for validation. Red rectangles highlight regions where segmentation performance is improved.